I need some help with some algebra word problems. I can't quite figure out the answers.
If you could help me i would greatly appreciate it.
The difference between the width and the length of a field is 12 m. Find its dimensions if a fence 160 m in length encloses it.
The length of a rectangle is found by doubling thw width and subtracting 3 m. Find its dimensions if the perimeter is 84 m.

Answer by checkley75(3666)

2L+2W=160 & W-L=12 THUS W=L+12
2L+2(L+12)=160
2L+2L+24=160
4L=160-24
4L=136
L=136/4
L=34 M FOR THE LENGTH.
W-34=12
W=12+34
W=46 M FOR THE WIDTH.
PROOF
2*34+2*46=160
68+92=160
160=160
------------------------------------------------
2W+2(2W-3)=84
2W+4W-6=84
6W=84+6
6W=90
W=90/6
W=15 M. FOR THE WIDTH
L=2*15-3
L=30-3
L=27 M. FOR THE LENGTH.
PROOF
2*15+2*27=84
30+54=84
84=84